OpenStudy (kawaiiiimya):

Please help An anthropologist finds bone that her instruments measure it as 0.824% of the amount of Carbon-14 the bones would have contained when the person was alive. How long ago did the person die? The half life of carbon 14 is 5,730 years.

OpenStudy (phi):

use \[ \text{ fraction left } = \left(\frac{1}{2}\right)^\frac{t}{5730} \]

OpenStudy (kawaiiiimya):

so (1/2) ^ 0.824/5730

OpenStudy (phi):

the 0.824% is how much is left of 100% or as a fraction (divide by 100 to get rid of the % sign) 0.00824 is how much is left. what we need to find is the time "t": \[ 0.00824 = \left(\frac{1}{2}\right)^\frac{t}{5730}\]

OpenStudy (kawaiiiimya):

LN 0.00824=(T/5730) LN 1/2

OpenStudy (phi):

yes, that looks good

OpenStudy (phi):

people solve these types of problems with logarithms. if you "take the log" of both sides you would get \[ \log(0.00824) = \log\left( \left(\frac{1}{2}\right)^\frac{t}{5730}\right) \]

OpenStudy (kawaiiiimya):

-4.798754935=(T/57300)(-0.693147181)

OpenStudy (kawaiiiimya):

6.923139943=(T/57300)

OpenStudy (phi):

ok, but it is 5730 (not 57300) so T= 5730 * 6.923139943= 39,669.59 years or about T= 39,700 years